OpenStudy (anonymous):

y=x-x^3.. solve for y y=x-x^3.. solve for y @Mathematics

OpenStudy (anonymous):

??

OpenStudy (anonymous):

sorry solve for x

jhonyy9 (jhonyy9):

y=x-x3 this is all ?

OpenStudy (anonymous):

yes. i want x=?

OpenStudy (lalaly):

:S http://www.wolframalpha.com/input/?i=y%3Dx-x^3+for+x

OpenStudy (anonymous):

complex isn.t it? that why i gave up on it!!

OpenStudy (lalaly):

Lol thats why i used wolfram

OpenStudy (anonymous):

y = x - x^3 y = x(1 - x^2) y = x(1^2 - x^2) y = x(1 + x)(1 - x) Stuck...

OpenStudy (anonymous):

ohh wow

OpenStudy (anonymous):

i give up too

jhonyy9 (jhonyy9):

why ?